If you pay 1.5% per month on the out standard balance of your master card bill what effective annual rate you being charged?

1
Expert's answer
2021-11-03T16:42:05-0400

Monthly: i = r/M

= 1.5%/12

=0.125%

Effective annual rate i a=(1 + r/M)M -1

Therefore; i a =(1 + 0.000125)12 - 1

=1.001501 - 1

=0.001501

=0.15%